OpenStudy (anonymous):

If the probability of success during a single event of a geometric experiment is 0.34, what is the probability of success on the 6th event? Round your answer to the nearest tenth of a percent. A. 4.3% B. 6.5% C. 2.8% D. 9.8%

OpenStudy (agent0smith):

If the p(success) = 0.34, then p(failure) = 1-0.34 = 0.66 I'm assuming you want it to fail the first five times, and succeed on the sixth: \[\Large 0.66^5 \times 0.34\]